Prev Next

SW&IT Meeting Minutes 2019-11-25

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191118
LR20191118 - Bec, Matthieu 40.0
LR20191118 - Holmberg, Carl 32.0
LR20191118 - Colson, Andrew 40.0
LR20191118 - Edmeades, David 0.0
LR20191118 - Sola, Igor 0.0
LR20191118 - Kubanek, Petr 24.0
LR20191118 - Hooper, Stephen 64.0
LR20191118 - Zhang, Yang 40.0
LR20191118 - Zhang, Xianyu 40.0

News from management

  • Payroll Processing Schedule: Time Sheets must be submitted and approved no later than 5:00 pm on Wednesday, November 27, 2019

Previous Action Items

Open topics

  • What is the state/report on the 64bit machine migration?
    • what goals have been reached?
    • who/what are we waiting for?
    • next steps?
  • [netmgrs] Please Share: Changes to the Microsoft Campus Agreement for Faculty and Staff

Team reports

XZ

Last week:
  • 4D trigger test
  • IDL license dicussion/test with IT
  • Reponse to SOUL FLAO-like operability
  • SOUL lab support
  • AdSec dump cause analyze
  • 4D work support
  • Start setup the mountain anallyze VM

This week:
  • Setup the mountain anallyze VM
  • 4D test support
  • AO telemetry document
  • SOUL lab setup support?

DE

  • Isilon wrangling; likely to defer pending deeper understanding of our needs, direction, and the system itself
  • PEPSI server order
  • VMWare support re-up finally ordered
  • 14TB disk test success; ordered remainder of disks for mountain Synology volume
  • Sorted Facilities on projector support panel move
  • IDL license investigations
    • Settle on node-locked for those machines that requrie it

YZ

  • LBC/MODS:
    • Created Github repo for LBC OB GUI code (together with related documents/codes).
    • Reviewed remote room machine operations regarding LBC and MODS with Olga.
    • Will work on future 64-bit obs machine setup for LBC and MODS. First example will be the TCSSendWavefront.c program.
    • Needs access to a Fedora30 VM?
    • Did some checking on older MODS files with Olga regarding IT#7927 (MODS image corrupted during transfer). Discovered more corrupted files.
    • Possible MODS software handover discussion?
  • AO-Cravat:
    • Met with Al Conrad and Steve Ertel on Friday (mainly to discuss Strehl calculation).
    • Discussed briefly with Shane regarding OSCO image categorization.
    • Worked with Steve on a set of real OSCO images and analyzesed how his code would work (confirmation of correct Strehl number is still pending, Shane is also supposed to provide Strehl using a different method on the same dataset).
    • Due to new agreement on OSCO image metadata, will work on rebuilding a database to track OSCO image and sequence (and provide the results to Al, Steve, Shane for reference).
    • Needs to rethink about some AO-Cravat interface (Al also has some new requirements as well).

PK

  • NewGCS wiki page
  • azcam-sim git & Docker integration (see README.md how to run the simulator)
  • azcam-sim image generation using photutils (which really rocks for astronomy image processing)
  • sorted out GCS configuration nightmare - getting rid of the TCSConfig in favour of Configuration, TCSConfig is used only for GCS
  • AZCam class able to connect from a config file, read in intialization

MB

  • OC, WPC, Management
  • Met SHARK VIS team, update all SHARKs + Arcetri on s/w interface
  • Work on a new document for CV detailing s/w requirements for new instruments
  • Posting for new position circulated and accepted, GG next for posting
  • Meeting on shutdown/restart and documenting procedure
  • Meeting on sciarc, Al's 3 tables and indirection concept that has potential to solve few problems
  • Look into few ITs (PMC packets, MODS striped images after transfer)
  • Meet and discuss foresesable "big items" with IT
next
  • Obs WS presentaton to sciops if time allows
  • Power restart document
  • DMS hashed urls, ordering issue at startup

IS

Last week:
  • miniQ
    • finish report part, routing and history navigation.
  • PIT
    • worked on ways to improve and automatize proposals after period closes.
    • worked on implementing a resubmit option to the PIT so there are no or to limit duplicates
This week
  • miniQ
    • clean up and make ingesting or setting up new run, easier (only from backend)
  • PIT
    • ongoing work on resubmission logic
  • LUCI
    • Need to do some code repository cleaning

AC

  • PMS - waiting for support from LASAIR
  • OAC - need to update git (IT)
  • ARGOS     
    • doc handover and prep
    • Jenkins RPM build - successfully builds TaN revision 15543 but later fails: "Build step 'Execute shell' marked build as failure"

CH

Last week
  • ALH: Clarified LBTI alerting reqs and began to sketch out an implementation based on an INDI module.
  • ARGOS: Reviewed design docs.
  • Supportscripts: Collected data on DS9 crashes. One appears to be a libxcb bug.
  • TMS: Added a test to directly test up/down state of TMS host.
  • Safety training.

This week
  • ALH: Prototype an LBTI agent.
  • ARGOS: Review design docs.
  • Supportscripts: Diagnose DS9 crashes during multiple script runs. Look at implementing a DS9 process tracking and restart function.
  • TMS: Hoping for opportunity to test status script/ALH integration.

SH

  • iLocater USB, UPS.
  • robs setup
    • robs32 installed.
    • robs installed (needs home directories, passwords).
    • Started slides
    • CentOS 8 Stream evaluated.
  • VM licensing
  • Olga VPN

next:
  • PEPSI keys
  • Olga IDL
  • ovirt on dms
  • startup docker registry
Topic revision: r16 - 25 Nov 2019, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback